OpenStudy (anonymous):

Estimate 8 and one-third times 2 and four-fifths using any method you like. Then find the actual product. Use your estimate to show that your answer is reasonable.

OpenStudy (anonymous):

ill be back in a min hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Answer: = 25/3 × 14/5 = (25 × 14) / (3 × 5) = 350/15 = 70/3 = 23 1/3

OpenStudy (anonymous):

if your gonna estimate it your answer would be 20 1/2

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

it would be 23 by itself

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np